# Prune Redundancy, Preserve Essence: Vision Token Compression in VLMs via Synergistic Importance-Diversity

### 10 Mar 2026 v1

@@ 3.3 Intra-Group Redundancy Removal via NMS
We then set the NMS threshold as $\tau=\lambda\cdot\rho$, where $\lambda$ is a scaling factor determined by the global token budget $N$. In our experiments, we empirically set $\lambda=\frac{N}{32}$, which consistently worked well across different compression settings. This adaptive threshold encourages stronger suppression for more redundant images.
After performing NMS for all groups, we obtain a collection of filtered groups $\{\widetilde{G}_{1},\ldots,\widetilde{G}_{K}\}$. To match a global token budget $N$, we allocate group-wise quotas $\{n_{1},\dots,n_{K}\}$ such that $\sum_{k=1}^{K}n_{k}=N$, where $n_{k}$ is calculated by rounding $\frac{|\widetilde{G}_{k}|}{\sum_{j}|\widetilde{G}_{j}|}\cdot N$ to the nearest integer.
@@ 3.4 Information-Aware Dynamic Compression Ratio Across Images
To address this limitation, we propose an information-aware dynamic compression ratio strategy that automatically adjusts the retained token budget $N$ according to each image’s information content. Building upon the global redundancy measure $\rho$ from Eq. 4, we compute an image information score:
@@ 3.5 Theoretical Overview of PSCA–NMS Pruning
Our pruning objective is to retain a token subset $S^{\prime}$ of fixed size $N$ that maximizes the effective information necessary for VLM. Formally, for a token set $S$, its informativeness is defined as:
@@ 4.3 Ablation Study
Ablation on Token Group Counts K. We study how the number of token groups ($K$) affects PruneSID performance under different total retained token counts ($N\in\{64,128,192\}$). For each $N$, we sweep the number of groups $K$ from 8 to 64. As shown in Tab. 6, performance exhibits a bell-shaped trend: too few groups reduce the granularity of redundancy modeling, while too many groups lead to overly small group sizes and unstable pruning. The optimal settings align with moderate values of $K=\frac{N}{4}$. These results validate our heuristic choice of increasing $K$ proportionally with $N$, balancing diversity and intra-group competition.
@@ A.1.1 Detailed Theoretical Analysis of PSCA–NMS Pruning
Given a retained subset $S^{\prime}$ of fixed size $N$, the informativeness can be expanded by the Inclusion–Exclusion Principle:
@@ A.3.3  Ablation Studies on Sensitivity to NMS Threshold Scaling Factor $\lambda$
This section presents a comprehensive ablation study to analyze the sensitivity of our method to the NMS threshold scaling factor $\lambda$, and its impact on both model performance and the distribution of retained tokens across semantic groups. As background, our NMS threshold is defined as $\tau=\lambda\cdot\rho$ (where $\rho$ denotes the redundancy metric between tokens), and $\lambda$ is a scaling factor adaptively determined by the global token budget $N$ (with a default setting of $\lambda=\frac{N}{32}$ in our base model). We design two complementary experiments to evaluate the role of $\alpha$ with $\lambda=\frac{N}{\alpha}$:

### 11 Mar 2026 v2

@@ 3.3 Intra-Group Redundancy Removal via NMS
We then set the NMS threshold as $\tau=\lambda\cdot\rho$, where $\lambda$ is a scaling factor determined by the global token budget NN. In our experiments, we empirically set $\lambda=\frac{N}{32}$, which consistently worked well across different compression settings. This adaptive threshold encourages stronger suppression for more redundant images.
After performing NMS for all groups, we obtain a collection of filtered groups $\{\widetilde{G}_{1},\ldots,\widetilde{G}_{K}\}$. To match a global token budget NN, we allocate group-wise quotas $\{n_{1},\dots,n_{K}\}$ such that $\sum_{k=1}^{K}n_{k}=N$, where $n_{k}$ is calculated by rounding $\frac{|\widetilde{G}_{k}|}{\sum_{j}|\widetilde{G}_{j}|}\cdot N$ to the nearest integer.
@@ 3.4 Information-Aware Dynamic Compression Ratio Across Images
To address this limitation, we propose an information-aware dynamic compression ratio strategy that automatically adjusts the retained token budget NN according to each image’s information content. Building upon the global redundancy measure $\rho$ from Eq. 4, we compute an image information score:
@@ 3.5 Theoretical Overview of PSCA–NMS Pruning
Our pruning objective is to retain a token subset $S^{\prime}$ of fixed size NN that maximizes the effective information necessary for VLM. Formally, for a token set $S$, its informativeness is defined as:
@@ 4.3 Ablation Study
Ablation on Token Group Counts K. We study how the number of token groups ($K$) affects PruneSID performance under different total retained token counts ($N\in\{64,128,192\}$). For each NN, we sweep the number of groups $K$ from 8 to 64. As shown in Tab. 6, performance exhibits a bell-shaped trend: too few groups reduce the granularity of redundancy modeling, while too many groups lead to overly small group sizes and unstable pruning. The optimal settings align with moderate values of $K=\frac{N}{4}$. These results validate our heuristic choice of increasing $K$ proportionally with NN, balancing diversity and intra-group competition.
@@ A.1.1 Detailed Theoretical Analysis of PSCA–NMS Pruning
Given a retained subset $S^{\prime}$ of fixed size NN, the informativeness can be expanded by the Inclusion–Exclusion Principle:
@@ A.3.3  Ablation Studies on Sensitivity to NMS Threshold Scaling Factor $\lambda$
This section presents a comprehensive ablation study to analyze the sensitivity of our method to the NMS threshold scaling factor $\lambda$, and its impact on both model performance and the distribution of retained tokens across semantic groups. As background, our NMS threshold is defined as $\tau=\lambda\cdot\rho$ (where $\rho$ denotes the redundancy metric between tokens), and $\lambda$ is a scaling factor adaptively determined by the global token budget NN (with a default setting of $\lambda=\frac{N}{32}$ in our base model). We design two complementary experiments to evaluate the role of $\alpha$ with $\lambda=\frac{N}{\alpha}$: